OpenStudy (anonymous):

could you help me solve 2h+b=132.26 h+2b=141.58

OpenStudy (anonymous):

2h+b=132.26 h+2b=141.58 here multiply 2, to eliminate the h part so that b part remain will give you 2h+4b=283.86 now subtract them from the first equation 2h+b=132.26 -(2h+4b=283.16) ---------------- 0- 3b =150.9 b=150.9/(-3)=50.3 answer

OpenStudy (anonymous):

http://www.myalgebra.com/default.aspx try plugging this equation into this website